OpenStudy (anonymous):

A=1/2(B+b)--------------A=258 B=20 b=23 FIND THE AREA OF A TRAPEZOID.

OpenStudy (lgbasallote):

didn't you just say A = 258?

OpenStudy (lgbasallote):

by the way...your formula for area of a trapezoid is wrong. Just saying.

OpenStudy (anonymous):

YES A =258

OpenStudy (lgbasallote):

and A means area right?

OpenStudy (anonymous):

ignore the words, just focus on the solving part lol :)

OpenStudy (lgbasallote):

i can't really ignore the words because you can't solve it without the right words

OpenStudy (anonymous):

its giving you the #s to plug in but I dont know how to solve it. Thats the formula its giving me . the words are just say to find the area. NOT to find A area.

OpenStudy (lgbasallote):

what i was saying is that you said A = 1/2 (B + b) and then you said A = 258 so im asking why are you looking for the area when it is given that A = 258 and i assume A means area

OpenStudy (lgbasallote):

and like i said before...your formula is wrong. it should be something like \[A = \frac{h(B+b)}2\]

OpenStudy (lgbasallote):

where h = height

OpenStudy (lgbasallote):

.....is what im saying making sense to you?

OpenStudy (anonymous):

thats how the problem is in my book